The theory predicts that the population of beans in the four groups A,B, C and D

should be 9:3:3:1. In an experiment among 1600 beans,the number in the four

groups was 882,313,287 and 118. Do theexperimental results support the survey


1
Expert's answer
2021-06-04T06:23:23-0400

"\chi^2" test

"H_0:" the experiment results support the theory

"H_a:" the experiment results do not support the theory


"\chi^2=\sum \frac{(O_i-E_i)^2}{E_i}=4.72"

Degrees of freedom:

"df=4-1=3"

From table:

"\chi^2_{0.05}=7.815"

Since the calculated value of "\chi^2"  is less than the table value, the null hypothesis may be accepted

with the significance level "\alpha=5\ \%"